A car standing from rest accelerates uniformly and attends a speed of 60ms-1 in 30 seconds. It then slows down uniformly until it comes to rest in the next 20 seconds. Sketch the velocity time graph for the motion of the car
1
Expert's answer
2019-10-31T11:18:39-0400


Acceleration during the first 30 seconds is "a_1 = \frac{60m/s}{30s} = 2 \frac{m}{s^2}" and during the next 20 seconds is "a_2 = \frac{-60m/s}{20s} = -3 \frac{m}{s^2}".